# Unlearning via Sparse Representations

### v1

@@ 4 Unlearning via Sparse Representations
Learning a Discrete Key Value Bottleneck.
A Discrete Key Value Bottleneck (DKVB) model is first trained on the given dataset using the standard negative log-likelihood (cross-entropy loss) training objective for multi-class classification. We use a non-parametric average pooling decoder and a CLIP (Radford et al., 2021) pre-trained ViT-B/32 (Dosovitskiy et al., 2020) as the backbone in all our experiments involving the DKVB. Then we proceed to unlearn a specific subset of data from these models. Before training with the classification objective, we do a key initialization for the DKVB models on the same dataset.
@@
Inference for Unlearning.
We propose to achieve unlearning in DKVB models by excluding key-value pairs from the bottleneck such that they cannot be selected again. This masking is done by setting the quantization distance of the selected keys to ‘infinity’. Figure 1 (center and right column) shows an overview of the proposed methods. More specifically, we experiment with two methods, Unlearning via Activations and Unlearning via Examples, described as follows.
@@
Unlearning via Activations and Unlearning via Examples are two different ways of achieving a common objective: to exclude a subset of activations corresponding to the forget set. However, using one approach over the other may be more practical or even necessary, depending on the task at hand. For selective unlearning where the goal is to forget specific examples, Unlearning via Examples would be necessary. In both the above approaches, we do not do any form of retraining or fine-tuning. Hence both these approaches require negligible additional compute.
@@ 5 Experiments and Results
The goal of our experiments is three-fold. First, we validate that we can zero-shot unlearn information via the Unlearning via Activations and Unlearning via Examples methods in models with a DKVB (Section 5.2). Second, we show that the proposed method is competitive with SCRUB (Section 5.3). Third, we show that the proposed method is equally competitive in the less constrained setting where we allow gradient-based re-training as is commonly required by previous methods (Section 5.4). Before presenting these results, we describe our experimental setup.
@@ 5.1 Experimental Setup
Benchmark datasets
We validate the proposed methods using experiments across three base datasets: CIFAR-10 with 10 distinct classes, CIFAR-100 (Krizhevsky et al., 2009) with 100 distinct classes and LACUNA-100 (Golatkar et al., 2020a) with 100 distinct classes. LACUNA-100 is derived from VGG-Faces (Cao et al., 2018) by sampling 100 different celebrities and sampling 500 images per celebrity, out of which 400 are used as training data and the rest are used as test images.
Models
On the aforementioned three datasets we study the following types of model architectures:
@@
    Backbone + Discrete Key-Value Bottleneck (Ours): Overall, this architecture consists of three components: 1) the frozen pre-trained backbone 2) the Discrete Key-Value Bottleneck (DKVB) and 3) a decoder, as shown in Figure 1. For the DKVB, we use 256 codebooks, with 4096 key-value pairs per codebook (approximately 1M pairs overall) following Träuble et al. (2023).
@@
    Backbone + Linear Layer (Baseline): As a baseline, we replace the Discrete Key Value bottleneck and the decoder in the above model architecture with a linear layer. Thus, the two components of this model are 1) a frozen pre-trained backbone and 2) a linear layer. This model will be used for the SCRUB baseline method.
@@
In each model, we use a pre-trained frozen CLIP (Radford et al., 2021) ViT-B/32 as our encoder backbone. We refer to the appendix for additional implementation details.
Training the Base Models
@@
We then train both model architectures on the full training sets of each dataset. Since the backbone is frozen, for the baseline, only the weights of the linear layer are tuned during both initial training (and later unlearning). Since we use only one linear layer, we do not do any sort of pre-training (beyond the backbone), unlike in previous works (Kurmanji et al., 2023; Golatkar et al., 2020a; b). Table 1 shows the performance of these trained models on the train and test splits of the complete datasets. Starting from these base models trained on the full datasets, we will validate the ability to unlearn previously learned knowledge.
@@
Table 1: Performance of the models on different sets of data after the initial training on the three datasets. We use two kinds of models: (a) models having a Discrete KV Bottleneck which are used for the proposed methods and (b) models where the DKVB and the decoder are replaced by a Linear Layer. These are used for the baseline. We wish to reduce the accuracy of these models on ${D}_{test}^{forget}$ to 0% while maintaining the accuracy on $\mathcal{D}_{test}^{retain}$.
Dataset $\mathcal{D}_{train}$ $\mathcal{D}_{train}^{retain}$ $\mathcal{D}_{train}^{forget}$ $\mathcal{D}_{test}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
CIFAR-10 100% 100% 100% 93.01% 92.61% 96.50%
CIFAR-100 99.98% 99.98% 100% 78.43% 78.24% 96.00%
LACUNA-100 98.09% 98.07% 100% 90.38% 90.28% 100%
@@
Dataset $\mathcal{D}_{train}$ $\mathcal{D}_{train}^{retain}$ $\mathcal{D}_{train}^{forget}$ $\mathcal{D}_{test}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
CIFAR-10 93.27% 92.82% 97.32% 93.02% 92.59% 96.90%
CIFAR-100 86.73% 86.61% 99.00% 78.53% 78.35% 96.00%
LACUNA-100 95.58% 95.53% 100% 90.68% 90.59% 100%

### v2

@@ 4 Unlearning via Sparse Representations
Learning a Discrete Key Value Bottleneck. A Discrete Key Value Bottleneck (DKVB) model is first trained on the given dataset using the standard negative log-likelihood (cross-entropy loss) training objective for multi-class classification. We use a non-parametric average pooling decoder and test the proposed approaches on two pretrained backbones: 1.) a CLIP (Radford et al., 2021) pre-trained ViT-B/32 (Dosovitskiy et al., 2020) and 2.) a ResNet-50 pretrained on ImageNet in a supervised fashion. Then we proceed to unlearn a specific subset of data from these models. Before training with the classification objective, we do a key initialization for the DKVB models on the same dataset.
@@
Inference for Unlearning. We propose to achieve unlearning in DKVB models by excluding key-value pairs from the bottleneck such that they cannot be selected again. Numerically, this masking is done by setting the quantization distance of the selected keys to ‘infinity’. Figure 1 (center and right column) shows an overview of the proposed methods. More specifically, we experiment with two methods, Unlearning via Activations and Unlearning via Examples, described as follows.
@@
Both approaches are two different ways of achieving a common objective: to exclude a subset of activations corresponding to the forget set. However, using one approach over the other may be more practical or even necessary, depending on the task at hand. In both the above approaches, we do not do any form of retraining or fine-tuning. The only computation which may be necessary is incurred during the inference stage for recording the key-value pairs which have been utilized for the forget set. Hence both approaches require negligible additional compute. Moreover, the requirement of access to original training data of the forget class can also be circumvented under appropriate assumptions, making the proposed approaches zero-shot unlearning methods.
@@ 5 Experiments and Results
The goal of our experiments is two-fold. First, we validate that proposed methods of the Unlearning via Activations and Unlearning via Examples in models with a DKVB (Section 5.2), and show that the proposed methods are competitive with the baselines (Section 5.2.1) in unlearning the forget class while incurring minimal damage to the performance of the models on the retain class. Second, we compare the compute efficiency of the proposed methods against that of the baselines. More specifically, we report the number of floating point operations (FLOPs) required during the procedure of unlearning. Before presenting these results, we describe our experimental setup.
@@ 5.1 Experimental Setup
Benchmark datasets We validate the proposed methods using experiments across four base datasets: CIFAR-10 with 10 distinct classes, CIFAR-100 (Krizhevsky et al., 2009) with 100 distinct classes, LACUNA-100 (Golatkar et al., 2020a) with 100 distinct classes and ImageNet-1k (Russakovsky et al., 2015) with 1000 distinct classes. LACUNA-100 is derived from VGG-Faces (Cao et al., 2018) by sampling 100 different celebrities and sampling 500 images per celebrity, out of which 400 are used as training data and the rest are used as test images.
Models On the aforementioned three datasets we study the following types of model architectures:
@@
    Backbone + Discrete Key-Value Bottleneck (Ours): Overall, this architecture consists of three components: 1) the frozen pre-trained backbone 2) the Discrete Key-Value Bottleneck (DKVB) and 3) a decoder, as shown in Figure 1. For the DKVB, we use 256 codebooks, with 4096 key-value pairs per codebook (approximately 1M pairs overall) as in Träuble et al. (2023).
@@
    Backbone + Linear Layer (Baseline): As a baseline, we replace the Discrete Key Value bottleneck and the decoder in the above model architecture with a linear layer. Thus, the two components of this model are 1) a frozen pre-trained backbone and 2) a linear layer. This model will be used for all the baseline methods.
@@
In each model, we use a pre-trained frozen CLIP (Radford et al., 2021) ViT-B/32 and ImageNet supervised pre-trained ResNet-50 as our encoder backbones. We refer the reader to the appendix for additional implementation details.
@@
Training the Base Models We then train both model architectures on the full training sets of each dataset. Since the backbone is frozen, for the baseline models, only the weights of the linear layer are tuned during initial training (and later unlearning). Since we use only one linear layer, we do not do any pre-training (beyond the backbone), unlike in previous works (Kurmanji et al., 2023; Golatkar et al., 2020a; b). Table 3 shows the performance of these trained models on the train and test splits of the complete datasets. Starting from these base models trained on the full datasets, we will validate the ability to unlearn previously learned knowledge.
Figure 2: Unlearning via Activations. Performance on the retain set test data vs. Performance on the forget set test data across various datasets for (a) CLIP pretrained ViT/B-32 in the top row (b) ImageNet pretrained ResNet-50 backbones in the bottom row as the value of $N_{a}$ is increased which is indicated by the color of the markers. The relative performance on the retain set test data as compared to the original models increases after unlearning in the case of CIFAR-10 and ImageNet-1k and drops for CIFAR-100 and LACUNA-100 in the case of ViT/B-32 and increases for all four datasets in the case of ResNet-50 (see Table 1).
Unlearning We aim to make the problem of unlearning as challenging as possible in order to fairly evaluate the proposed methods. Therefore, on each dataset we select the class that is best learned by the respective models with the Discrete Key Value Bottleneck trained previously, to be the forget class (see Appendix A.2 for further details).
Objective & Metrics We report our results on the test data of retain classes and forget class, i.e. $\mathcal{D}_{test}^{retain}$ and $\mathcal{D}_{test}^{forget}$. Further, in our experiments, we aim to achieve complete unlearning - achieving minimal accuracy on the forget set while incurring minimal damage to the performance on the retain set. While achieving complete unlearning may not always be desirable, such as in the case of Membership Inference Attacks (MIAs) the proposed methods can be easily extended to defend against MIAs (We refer to Appendix A.6 for further discussion on MIAs and the proposed methods). We report mean values across 5 random seeds in all cases.
For comparing the compute efficiency of different approaches, we report the approximate FLOPs (Floating Point Operations) required for the procedure of unlearning. The total number of FLOPs are calculated as number of FLOPs required during the forward passes + number of FLOPs required during the backward passes. We use the fvcore [^1] https://github.com/facebookresearch/fvcore library for computing the number FLOPs required during the forward passes. FLOPs required using backward passes are approximated as number of operations used for gradient computations + number of operations used for weight updates [^2] https://epochai.org/blog/backward-forward-FLOP-ratio. Since only the linear head weights are trainable, the number of computations required for calculating the gradients would be the same as the number of parameters in the linear layer. Further, since we use Adam optimizer, the number of operations required for the weight updates would be equal to 18 times the number of parameters.
To calculate the final number of FLOPs, we first calculate the FLOPs required for one example (one forward + backward pass) and then multiply them with the total number of examples and the total number of epochs. For SCRUB, the forward and backwards FLOPs are multiplied with different scalars depending on the msteps parameter.
5.2 Unlearning via the Discrete Key-Value Bottleneck
We will now discuss the results of unlearning via activations and examples, i.e. the two approaches proposed in Section 4 on all four benchmark datasets.
Figure 3: Unlearning via Examples. Performance on the retain set test data vs. Performance on the forget set test data across different datasets for (a) CLIP pretrained ViT/B-32 in the top row and (b) ImageNet pretrained ResNet-50 backbones in the bottom row as the value of $N_{e}$ is increased which is indicated by the color of the markers. The relative performance on the retain set test data as compared to the base model increases for CIFAR-10 and drops for all other datasets in the case of ViT/B-32, whereas it drops for CIFAR-10 and CIFAR-100 and increases for LACUNA-100 and ImageNet-1k in the case of ResNet-50 (see Table 1)
Unlearning via Activations. Unlearning via activations requires us to set the hyperparameter $N_{a}$, reflecting the top-$N_{a}$ most frequently activated key-value pairs which will be masked out after inference on the forget set. We therefore start by analyzing its role over a wide range of values to probe its choice and effect with $N_{a}=0$ being the limit without any unlearning. Figure 2 summarizes the unlearning and effect of $N_{a}$ on the retain vs forget test set. In the case of CIFAR-10 and a ViT/B-32 backbone, the initial accuracies on the retain and forget test set are 92.61% and 96.50% respectively. As $N_{a}$ increases, the forget class test accuracy decreases, slowly for small $N_{a}$ and rapidly for larger $N_{a}$. The model reaches random accuracy (i.e. 10% for CIFAR-10) on the forget class test data at $N_{a}=150000$. At this point the retain set test accuracy is 92.97%. The model unlearns the forget class completely between $N_{a}=170,000$ (0.4%) and $N_{a}=180,000$ (0%). At this point, the retain set test accuracy is 92.94%, which is almost identical to the initial accuracy. Further increasing $N_{a}$ up to $N_{a}=200,000$, i.e. about 20% of all key-value pairs, leads to an additional increase in retain set test accuracy to 93%. On the contrary, in the case of CIFAR-10 and a ResNet-50 backbone, the decrease in the forget set test accuracy is rapid for small $N_{a}$ and it slow for higher values of $N_{a}$. Complete unlearning in this case happens between $N_{a}$ = 160000 (0.1%) and $N_{a}$ = 190000 (0%). These differences in trends can be attributed to how the information is factorized among the representations. For eg. there is a steep decline in the forget set test accuracy between $N_{a}$ = 50000 (41%) and $N_{a}$ = 55000 (13%) in the case of LACUNA-100 and ResNet-50 backbone (Figure 2). This behavior may be attributed to the presence of high information but less frequently selected key-value pairs between the two values of $N_{a}$. Nevertheless, as can be seen from the equivalent analysis on the CIFAR-100, LACUNA-100 and ImageNet-1k models in Figure 2, the same trend of maintaining the initial retain accuracy while minimizing the forget accuracy up to a minimum holds across all four datasets and both backbones validating its meaningful unlearning capability.
Unlearning via Examples. For the second method—unlearning via examples—$N_{e}$ examples are sampled randomly from the training data of the forget class, and subsequently used for unlearning by the mechanism described in Section 4. Similar to before, we aim to assess the effect on the choice of $N_{e}$ over a wide range for each dataset, including $N_{e}=0$ being the limit without any unlearning. Figure 3 summarizes the unlearning and effect of $N_{e}$ on the retain vs. forget test set. We again begin by focusing on the results with CIFAR-10 and ViT/B-32 backbone. Here, the forget set $\mathcal{D}_{train}^{forget}$ contains 5000 examples. We start off with retain set and forget set test accuracies of 92.61% and 96.50% respectively. Similar to the previous approach – unlearning via activations – the test accuracy on the forget set decreases with increasing $N_{e}$. The accuracy on the retain test set, on the other hand, increases monotonically, although only slightly overall. The model achieves random accuracy on the forget class around $N_{e}=2500$. The accuracy on retain set test data is at just under 93% at this transition. Finally, the accuracy on the forget set drops to 0% (i.e. complete unlearning) between $N_{e}=3000$ and $N_{e}=3400$ with a retain set test accuracy of just above 93% at $N_{e}=3400$. Further increasing $N_{e}$ does not affect the retain set test accuracy notably. Similarly to the case of unlearning via activations, the forget set test performance decreases rapidly at first and then slowly with $N_{e}$ in the case of CIFAR-10 with a ResNet-50 backbone. The retain set test accuracy increases at first and then decreases, albeit marginally. An equivalent analysis on the CIFAR-100, Lacuna-100 and ImageNet-1k models in Figure 3 exhibits a similar behavior of successful minimization of the forget accuracy up to a minimum while roughly maintaining the retain set test accuracy, validating unlearning via examples as another option for unlearning using discrete key-value bottlenecks.
Summary. Both methods, Unlearning via Activations and Unlearning via Examples, successfully demonstrate unlearning of the forget class while having a negligible effect on the models’ performance on the retain set. Importantly, this is achieved without any form of training, retraining, or fine-tuning as is usually required by other methods. The retain set test accuracy remains more or less constant for all four datasets except for a few minor fluctuations. This is a result of the fact that due to localized and context-dependent sparse updates during the initial training of the model, discrete key-representations corresponding to different classes in the dataset are well separated from each other, an important prerequisite discussed in Träuble et al. (2023). Hence, all the information about a class can be unlearned by forgetting only a subset of the forget class training data in the case of Unlearning via Examples, making it very data-efficient.
Table 1: Comparison between the proposed methods and the baseline across CIFAR-10, CIFAR-100, LACUNA-100 and ImageNet-1k datasets and CLIP pretrained ViT/B-32 and ImageNet pretrained ResNet-50 backbones. We compare the relative change in performance on the retain and forget set test data relative to the originally trained models. The proposed methods are able to unlearn the forget sets completely in all cases while causing minimal changes in the performance of the models on the retain set test data.
CIFAR-10 CIFAR-100 LACUNA-100 ImageNet-1k
Backbone Method $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
ViT/B-32 DKVB via Activations (sec 5.2) 0.36% -100% -0.20% -100% -0.17% -100% 0.15% -100%
DKVB via Examples (sec 5.2) 0.45% -100% -0.36% -100% -0.09% -100% -0.03% -100%
Linear Layer + SCRUB 1.62% -100% -0.91% -100% -1.10% -100% 7.31% -100%
Linear Layer + Finetuning 1.94% -100% -1.91% -98.33% -2.21% -100% 0.88% -100%
Linear Layer + Retraining 1.82% -100% -0.39% -100% -2.03% -100% 5.16% -100%
Linear Layer + NegGrad+ 0.49% -100% -0.63% -100% -1.34% -100% 2.45% -100%
ResNet-50 DKVB via Activations (sec 5.2) 0.04% -100% 0.26% -100% 0.21% -100% 0.04% -100%
DKVB via Examples (sec 5.2) -0.07% -100% -0.34% -100% 0.17% -100% 0.04% -100%
Linear Layer + SCRUB -0.07% -99.67% -0.94% -98.79% -0.26% -99.67% 0.74% -100%
Linear Layer + Finetuning 0.48% -100% -0.46% -99.99% -2.96% -100% -2.25% -100%
Linear Layer + Retraining 3.06% -100% 1.76% -100% 1.15% -100% -1.14% -100%
Linear Layer + NegGrad+ 2.13% -100% -0.85% -100% 6.73% -100% -0.85% -100%
5.2.1 Comparison with Baselines
We now compare the results of both the proposed methods, which require Backbone + DKVB models against several baseline methods, which are optimized for models without such a bottleneck. For this, we will use the Backbone + Linear Layer models described in 5.1. On these models, we run SCRUB (Kurmanji et al., 2023), finetuning - finetuning the model to be unlearnt on the retain set, retraining - training the model from scratch on the retain set only and NegGrad+ (Kurmanji et al., 2023) and compare the performance changes on the forget and retain classes against the performance changes after unlearning with the two proposed methods.Table 1 shows the comparison between the two previously reported methods and the baselines. We can see that one of the two proposed approaches always results in the least change in the performance of the base model on the retain classes, while at the same time achieving complete unlearning of the forget class. The baselines on the other hand, occasionally fail to achieve complete unlearning. Finally, it is important to re-emphasize that the proposed methods achieve the shown performance without requiring any additional gradient-based training for unlearning. In the case of baselines, we stop the unlearning procedure when the forget set is completely unlearned or the forget set test accuracy has converged with minimal damage to the performance on the retain set. Moreover, while we report results for the case of complete unlearning, the proposed methods can be easily used for achieving unlearning of the forget class to different extents by tuning the $N_{a}$ and $N_{e}$ hyperparameters. We refer to Appendix A.7.3 and A.7.2 for further training and implementation details.
5.3 Proposed Methods achieve Unlearning in a Compute Efficient Manner
In this section, we compare the proposed approaches against the baselines in terms of the amount of compute required in order to achieve complete unlearning. To facilitate this comparison, we report the number of FLOPs required for the unlearning procedure for each case. The FLOPs are calculated following the rules described in Section 2. Table 2 compared the FLOPs required for unlearning in each case.
Table 2: Comparison of FLOPs for various methods across CIFAR-10, CIFAR-100, LACUNA-100, and ImageNet-1k datasets using ViT/B-32 and ResNet-50 backbones. We report both forward and backward FLOPs for each method.
CIFAR-10 CIFAR-100 LACUNA-100 ImageNet-1k
Backbone Method Forward Backward Forward Backward Forward Backward Forward Backward
(TFLOPs) (GFLOPs) (TFLOPs) (GFLOPs) (TFLOPs) (GFLOPs) (TFLOPs) (GFLOPs)
ViT/B-32 DKVB via Activations (sec 5.2) 21.93 0 2.19 0 1.75 0 5.63 0
DKVB via Examples (sec 5.2) 14.91 0 0.75 0 1.40 0 2.90 0
Linear Layer + SCRUB 655.13 14.59 1316.83 293.30 527.60 117.51 39168.28 87230.98
Linear Layer + Finetuning 196.54 4.38 6485.87 1444.61 5188.69 1155.69 11179.75 24898.21
Linear Layer + Retraining 196.54 4.38 1080.98 240.77 864.78 192.61 5589.87 12449.11
Linear Layer + NegGrad+ 393.08 8.76 1729.56 385.23 1383.65 308.18 11179.75 24898.21
ResNet-50 DKVB via Activations (sec 5.2) 16.66 0 1.67 0 1.33 0 5.31 0
DKVB via Examples (sec 5.2) 7.33 0 0.93 0 1.33 0 3.74 0
Linear Layer + SCRUB 1498.74 87.55 1488.79 869.68 3697.00 2159.62 31873.28 299227.17
Linear Layer + Finetuning 1049.12 61.29 1648.66 963.07 131.89 77.05 5304.25 49796.43
Linear Layer + Retraining 659.47 385.23 1648.66 963.07 2242.18 1309.78 5304.25 49796.43
Linear Layer + NegGrad+ 329.73 192.61 10608.49 99592.85 263.79 154.09 5304.25 49796.43
We report the forward and backward FLOPs separately to highlight that the proposed approaches do not require any gradient based updates. Additionally, while the scale of backward FLOPs may seem insignificant against the forward FLOPs, it can easily blow up in cases where complex parametric decoders are used on top of the DKVB. In our experiments, the decoder is simply an average pooling layer. Nevertheless, we can see that the proposed approaches require significantly less forward FLOPs as compared to the baselines. This can be explained by the fact that the proposed approaches require only one forward pass through the models per example of the forget class training data, in order to cache the activations. The baseline methods on the other hand, require multiple forward passes, each corresponding to a single training epoch.
6 Limitations and Future Work
The proposed methods inherit the limitations of the DKVB (Träuble et al., 2023) such as the reliance of DKVB on pre-trained encoders which can extract meaningful shared representations and trade-offs in downstream performance due to the use of an information bottleneck. Extensions to the model may involve training sparse representations inducing discrete bottleneck end-to-end. Further, in our experiments, we consider the setting of class incremental learning where the forget set can be easily identified and isolated. However, this may not always be sufficint for a given task and more complicated approaches might be needed to identify the data that needs to be removed from the model. While the methods introduced in this work are currently not designed for selective unlearning as outlined in Appendix A.6, there are various directions for future adaptations. These directions include enhancing the forget set isolation, and addressing limitations related to information retention, for instance by further fine-tuning on the encoder backbone. Advancing towards even stronger unlearning forms involving discrete key value bottlenecks, represents a key direction for future work.
7 Conclusion
In this work, we proposed a new approach to unlearning that requires minimal computation in order to unlearn a subset of data. This approach is based on the use of a discrete architectural bottleneck which induces sparse representations. These sparse representations facilitate unlearning a subset of data from the model with minimal to no performance drop on the rest of the data. We focused on the setting of class unlearning and our experiments show that the proposed approach, while being at least 20$\times$ compute efficient, performs competitively with or in some cases better than a state-of-the-art approach which requires additional compute to perform unlearning. Consequently, excising the activated key-value pairs from the model is a highly effective means of unlearning the forget set without disrupting the retain set.
8 Acknowledgements
This research was enabled by compute resources provided by Mila (mila.quebec). VS would like to thank Aniket Didolkar and Moksh Jain for useful discussions, feedback on the initial versions of the paper and proof reading the paper. VS would also like to thank Nanda H Krishna and Moksh Jain for helping with the figures included in the paper and Thomas Jiralerspong for proof reading and giving feedback on the paper.
Appendix A Appendix
A.1 Initial Performances of the Models
We train both: the models with a DKVB and the baseline models (i.e. backbone + linear layer) to achieve similar performances on the test datasets in order to ensure a fair comparison. Note that, due to these the models are not necessarily trained to achieve the maximum possible performance on the datasets. Table 3 shows the initial performances of the originally trained models on different splits of the datasets.
@@ A.1 Initial Performances of the Models
Table 3: Performance of the models on different sets of data after the initial training on the four datasets. We use two kinds of models: (a) models having a Discrete KV Bottleneck which are used for the proposed methods and (b) models where the DKVB and the decoder are replaced by a Linear Layer. These are used for the baseline. We wish to reduce the accuracy of these models on ${D}_{test}^{forget}$ to 0% while maintaining the accuracy on $\mathcal{D}_{test}^{retain}$.
ViT-B/32 ResNet-50
Dataset $\mathcal{D}_{train}$ $\mathcal{D}_{train}^{retain}$ $\mathcal{D}_{train}^{forget}$ $\mathcal{D}_{test}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{train}$ $\mathcal{D}_{train}^{retain}$ $\mathcal{D}_{train}^{forget}$ $\mathcal{D}_{test}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
CIFAR-10 100% 100% 100% 93.01% 92.61% 96.50% 100% 100% 100% 82.94% 82.04% 91.00%
CIFAR-100 99.98% 99.98% 100% 78.43% 78.24% 96.00% 99.98% 99.98% 100% 62.11% 61.81% 92.00%
LACUNA-100 98.09% 98.07% 100% 90.38% 90.28% 100% 98.35% 98.34% 100% 65.53% 65.24% 94.00%
ImageNet-1k 99.53% 99.53% 100% 68.24% 68.22% 92.00% 99.37% 99.36% 100% 76.44% 76.41% 100%
@@
ViT-B/32 ResNet-50
Dataset $\mathcal{D}_{train}$ $\mathcal{D}_{train}^{retain}$ $\mathcal{D}_{train}^{forget}$ $\mathcal{D}_{test}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{train}$ $\mathcal{D}_{train}^{retain}$ $\mathcal{D}_{train}^{forget}$ $\mathcal{D}_{test}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
CIFAR-10 93.27% 92.82% 97.32% 93.02% 92.59% 96.90% 83.80% 83.42% 87.20% 82.04% 81.74% 84.70%
CIFAR-100 86.73% 86.61% 99.00% 78.53% 78.35% 96.00% 78.02% 77.82% 98.20% 62.69% 62.41% 90.00%
LACUNA-100 95.58% 95.53% 100% 90.68% 90.59% 100% 86.48% 84.53% 99.25% 65.40% 65.10% 95.00%
ImageNet-1k 73.13% 73.11% 96.15% 68.29% 68.26% 92.00% 97.44% 97.43% 100% 76.77% 76.75% 100%